Suppose, T in Hom(V,W) then prove that Nullspace of T is subspace of V
1
Expert's answer
2020-06-14T18:14:06-0400

Let "v_1"  and "v_2"  belong to the nullspace of "T" . Then "T(v_1) = 0"  and "T(v_2) = 0"  by the definition of nullspace. Then "T(v_1+v_2) = T(v_1) + T(v_2) = 0+0 = 0"  because "T"  is linear, and "v_1+v_2"  belongs to the nullspace of "T" . Therefore, the nullspace of "T"  is closed under addition.


Since "T(0) = 0"  because "T"  is linear, the nullspace of "T"  contains .


Let "k"  be a scalar and "v"  belong to the nullspace of "T" . Then "T(v) = 0"  by the definition of nullspace. Then "T(kv) = k\cdot T(v) = k\cdot 0 = 0"  because "T"  is linear, and "kv"  belongs to the nullspace of "T" . Therefore, the nullspace of "T"  is closed under multiplication by a scalar.


By the definition of subspace, the nullspace of "T"  is a subspace of "V" .